Main Logbooks page LST commissioning Shift checklists Safe state checklists Maintenance activities LST2-4 Commissioning Old logbooks
GENERAL CAMERA DRIVE ENERGY AMC
  LST1 Drive ELOG  ELOG logo
Message ID: 251     Entry time: Thu Apr 18 21:21:25 2019
Author: Drive Team- Ino&Thierry&Alessandro 
Type: Telescope Operations 
Category: Hardware 
Subject: Elevation limit switch issue 
Status:  

The 2 swicthes are too far from the target at 56.48 elevation.  The sensor have been tuned in parkin position and the distance was correct (2mm) and at 56.48 is about 10mm. (impossible to tune correctly the sensor and the atrger has to be tuned)

ELOG V3.1.5-fc6679b